solve using any method
6x^2-x-2=0

Answer by mananth(16946)

6x^2-x-2=0
6x^2-4x+3x-2=0
2x(3x-2)+1(3x-2)=0
(3x-2)(2x+1)=0
x=2/3 OR x=-1/2
...
m.ananth@hotmail.ca